# A delayed interior area-to-height estimate for the Curve Shortening Flow

### v1

@@
Abstract
In [ST24a, ST24b], Topping & the author proposed the principle of delayed parabolic regularity for the Curve Shortening Flow; in [ST24a], they provided a handful of proper graphical situations in which their delayed regularity framework is valid. In this paper, we show that there holds an interior graphical estimate for the Curve Shortening Flow in the spirit of the proposed framework. More precisely, we show the following $\mathrm{L}^{1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate: If a smooth Graphical Curve Shortening Flow $u:(-1,1)\times[0,T)\mapsto\mathbb{R}$ starts from a function $u_{0}:=u(\,\cdot\,,0):(-1,1)\mapsto\mathbb{R}$ with $\mathrm{L}^{1}(\!(-1,1)\!)$–norm strictly less than $\pi T$, then after waiting for the magic time $t_{\star}:=\|u_{0}\|_{\mathrm{L}^{1}(\!(-1,1)\!)}/\pi$, the size $|u(0,t)|$ of $u(\,\cdot\,,t)$ at the origin at any time $t\in(t_{\star},T)$ is controlled purely in terms of $\|u_{0}\|_{\mathrm{L}^{1}(\!(-1,1)\!)}$ and $t-t_{\star}$. We apply our estimate to construct Graphical Curve Shortening Flows starting weakly from Radon measures without point masses.
@@ 1 Introduction
Following on from Huisken’s seminal paper ([Hui84]) on the MCF of smooth, closed and mean-convex hyper-surfaces in $\mathbb{R}^{n+1\geq 3}$, Gage & Hamilton tackled the question of the long-time behaviour of the CSF. In [GH86] they showed that the evolution $[0,T)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$ of a given convex smooth loop $\gamma_{0}=:\gamma(0)$ exists for a maximal time $T<\infty$ dictated only by the area of the region enclosed by $\gamma_{0}$. Moreover, they precisely described the behaviour of the evolution of the loop as time approaches $T$: the curve $\gamma(t)$ smoothly approaches a circle of radius ${\sqrt{2(T-t)}}$ in the sense of blow-ups as $t\nearrow T$. Shortly after, Grayson showed in [Gra87] that the evolution of an arbitrary smooth loop in the Euclidean plane becomes convex before encountering any singularities, at which point Gage & Hamilton’s work applies.
@@
Some of the pioneering works in this direction are [EH89, EH91] due to Ecker & Huisken. They showed that, given any [^2] Ecker & Huisken’s work considered the more general case of the MCF of graphical hyper-surfaces. graph $\gamma_{0}=\mathrm{graph}(u_{0})\subset\mathbb{R}^{2}$ of a locally Lipschitz function $u_{0}\in\mathrm{C}_{\mathrm{loc}}^{0,1}(\mathbb{R})$, there exists a CSF $(0,\infty)\ni t\mapsto\gamma(t)=\operatorname{graph}(u(t)\!)\subset\mathbb{R}^{2}$, defined for all time, which attains the initial datum $\gamma_{0}$ locally uniformly, and which remains graphical over the same line; we call this flow the the Ecker–Huisken flow of $u_{0}$. They achieved their existence theory by appealing directly to the Graphical Curve Shortening Flow (GCSF) equation, which the graphing function [^3] We write $u(t)=u(\,\cdot\,,t)$, and use throughout this paper the same convention for all functions of space-time. $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,T)\!)$ necessarily satisfies. A key step in their proof is to establish a priori estimates for the GCSF equation of smooth graphing functions, which propagate an initial local Lipschitz bound forward to the CSF, provided one moves in a little. Such bounds are stable under locally uniform limits, and thus allowed Ecker & Huisken to construct solutions via an approximation procedure.
Around the same time, Angenent developed his Intersection Principle (see [Ang88]) and used it in [Ang91] to construct solutions to the CSF starting from Lipschitz loops in the uniform sense. As in Ecker & Huisken’s work, a key step is to establish an a priori estimate for smooth loops which propagates a Lipschitz bound forward to the flow. Angenent introduced a new and remarkably geometric technique to establish his estimates, which has found much success in other situations—see e.g. [Ilm92, Clu04, NT05, BR24], and more. In particular, Topping & the author used Angenent’s technique in [ST24a] to establish sharp delayed estimates for the GCSF, which fit into the framework to be recalled in §1.1.
@@
The running theme in the works [EH89, EH91, Ang91] above is that a strong regularity condition is shown to be preserved by the CSF. In the past three decades, significant progress has been made in the graphical situation towards instantaneous estimates, which take a weak initial regularity constraint and deduce a strong regularity condition on the CSF for strictly positive times. Evans & Spruck established in [ES92] (see also [CM04]) that an $\mathrm{L}^{\infty}_{\mathrm{loc}}$–bound on an initial graphing function induces an instantaneous $\mathrm{L}_{\mathrm{loc}}^{\infty}$–bound on the gradient of the subsequent flow, provided one moves in a little. Their result was later extended, independently, by Clutterbuck and Nagase & Tonegawa in [Clu04] and [NT05] respectively; in particular, Clutterbuck used her generalisation to establish an existence theory for a family of parabolic equations which contains graphical instances of the CSF and MCF. For $\mathrm{L}^{p>1}_{\mathrm{loc}}$–bounds, Chou & Kwong established in [CK20] an a priori estimate which takes an initial $\mathrm{L}^{p>1}_{\mathrm{loc}}$–bound on a graphing function instantaneously to a $\mathrm{L}^{\infty}_{\mathrm{loc}}$–bound on the gradient of the subsequent flow, provided one moves in a little; they used their estimate to establish the existence of GCSFs starting from graphs of $\mathrm{L}^{p>1}_{\mathrm{loc}}(\mathbb{R})$–initial data. They also established a uniqueness result from $\mathrm{L}^{p>1}(\mathbb{R})$–initial data, building on Chou & Zhu’s uniqueness theory ([CZ98]); see also [DS23].
For the non-graphical case, see, for example, the works of Angenent ([Ang90]; loosely, initial curves with $\mathrm{L}^{p\gg 1}$–curvature) and Lauer ([Lau13]; finite-length Jordan curves).
One of the aims of this paper is to extend the existence theory to include, loosely, graphs of Radon measures without point masses. In some sense, the strategy we follow is familiar: we first establish an a priori estimate for smooth initial data which is stable under weak convergence of measures, and then use an approximation procedure to pass to a non-smooth limit. However, the a priori estimate we establish is pronouncedly unfamiliar in comparison to the previously-mentioned estimates [GH86, EH89, EH91, Ang90, Ang91, ES92, Lau13, CM04, Clu04, NT05, CK20, DS23] and, to the author’s best knowledge, all other estimates (except, of course, for [ST24a]) from the CSF–literature, which, recall, start to hold instantaneously. Our estimate is a delayed parabolic a priori estimate, and in our setting it is, in fact, impossible for the estimate to hold before a so-called magic time, which is determined completely by the ‘area’ under the initial graph. In particular, we apply our estimate slightly differently to the applications in, for example, [EH89, EH91, Ang90, Ang91, Lau13, Clu04, CK20].
@@ 1.1 A novel framework
In [ST24a, ST24b], Topping & the author proposed the following novel framework by which the CSF regularises a curve.
Given two families $[0,T)\ni t\mapsto\gamma^{-}(t)\subset\mathbb{R}^{2}$ and $[0,T)\ni t\mapsto\gamma^{+}(t)\subset\mathbb{R}^{2}$ of properly embedded curves in the Euclidean plane evolving under the Curve Shortening Flow, for which $\gamma^{-}(t)$ and $\gamma^{+}(t)$ are disjoint for each $t\in[0,T)$ and bound an evolving connected region of fixed area $\mathcal{A}$, define the magic time to be $t_{\star}:=\mathcal{A}/\pi$. Then, after flowing beyond the magic time $t_{\star}$, one expects to start being able to control the regularity of $\gamma^{+}(t)$ for times $t>t_{\star}$ in terms of $t-t_{\star}$, $\mathcal{A}$ and the regularity of $\gamma^{-}(t)$.
Several situations in which the framework above holds exactly as stated were provided in [ST24a]. Most relevant to this paper is the case when $t\mapsto\gamma^{-}(t)$ is the static flow of the $x$–axis:
@@
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times[0,\infty);\mathbb{R}_{\geq 0})$ be a positive Graphical Curve Shortening Flow starting from $u_{0}:=u(0)\in\mathrm{L}^{1}(\mathbb{R};\mathbb{R}_{\geq 0})$. Let
(1.2) $\mathcal{A}=\|u_{0}\|_{\mathrm{L}^{1}(\mathbb{R})}=\int_{x\in\mathbb{R}}u_{0}(x)\,\mathrm{d}x$
@@
(1.3) $t_{\star}=\frac{\mathcal{A}}{\pi}.$
@@
(1.4) $\|u(t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leq\sqrt{t}\cdot H\left(\frac{t}{\mathcal{A}}\right)$
@@
Recall that a GCSF is a function $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J\times(0,T)\!)$ which satisfies the quasi-linear parabolic equation
(GCSF) $\partial_{t}u=\frac{1}{1+(\partial_{x}u)^{2}}\partial^{2}_{x}u=\partial_{x}\!\left(\tan^{-1}(\partial_{x}u)\!\right),$
where $J\subseteq\mathbb{R}$ is an interval. A solution to (GCSF) induces a family of curves $(0,T)\ni t\mapsto\gamma(t)=\operatorname{graph}(u(t)\!)$ which solve the CSF equation
(CSF) $\left\langle\partial_{t}\gamma,\bm{n}\right\rangle=\kappa,$
@@
When an initial condition, say $u_{0}=:u(0)$ for (GCSF) or $\gamma_{0}=:\gamma(0)$ for (CSF), is imposed, the regularity of the initial datum determines the mode in which the flow converges backward in time. For example, in Chou & Kwong’s work [CK20], since $u_{0}\in\mathrm{L}^{p>1}_{\mathrm{loc}}(\mathbb{R})$, they insist that $u(t)\to u_{0}\in\mathrm{L}^{p>1}_{\mathrm{loc}}(\mathbb{R})$ as $t\searrow 0$. In the case that $\gamma_{0}$ is continuous, we write $\gamma:J\times[0,T)\mapsto\mathbb{R}^{2}$ or $[0,T)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$. In the case that $u_{0}\in\mathrm{C}^{0}_{\mathrm{loc}}(\mathbb{R})$, we write $u\in\mathrm{C}^{0}_{\mathrm{loc}}(\mathbb{R}\times[0,T)\!)$. And in the case that $u_{0}\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R})$, we write $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times[0,T)\!)$.
In [ST24a], Topping & the author also provided the following construction, which shows that the proposed magic time $t_{\star}:=\mathcal{A}/\pi$ is sharp:
@@
(1.7) $\sup_{n\in\mathbb{N}}v_{n}(0,s)=\infty\qquad\text{ and }\qquad\sup_{n\in\mathbb{N}}\|v_{n}(t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}<\infty.$
As an application of Theorem 1.1, Topping & the author established the existence of GCSFs starting from $\mathrm{L}^{1}(\mathbb{R})$–initial data. The existence result in this paper is a generalisation of the existence result of Topping & the author, but uses a slightly different method.
@@ 1.2 Main results
Throughout this paper, we use the notation $J_{r}(a)$ to mean the interval $a+(-r,r)\subset\mathbb{R}$ thought of as a spatial domain. We also abbreviate $J_{r}(0)$ to $J_{r}$. Any instance of a $J$ without the decoration as before will refer to an interval of unspecified length, centre or closed-ness. This is done purely to distinguish the spatial and temporal domains at the level of notation.
The crucial new estimate in this paper is
Theorem 1.8 (Delayed $\mathrm{L}^{1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$ estimate).
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a Graphical Curve Shortening Flow starting from $u_{0}:=u(0)\in\mathrm{L}^{1}(J_{1})$, and suppose that
(1.9) $t_{\star}:=\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}=\frac{1}{\pi}\int_{y\in J_{1}}|u_{0}(x)|\,\mathrm{d}x<T.$
Then, for all $\delta>0$ there exists a constant $C=C(\delta)<\infty$ such that the bound
(1.10) $|u(0,t)|\leq C+\frac{1}{2}\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}+\frac{\pi t}{2}$
hold for all $t\geq(1+\delta)\cdot t_{\star}$ (provided, of course, that $t<T$ too).
@@
Once a height bound has been established, Evans & Spruck’s estimate ([ES92]) implies that an interior gradient bound holds an instant later, at which point standard parabolic regularity theory (see e.g. [LSU88, Kry96]) can be used to deduce that higher order bounds hold after one more instant. For details, see [Sob24]. We shall use this argument in the proof of our existence result, which will be stated momentarily.
@@
As an application of Theorem 1.8, in §4 we relatively swiftly deduce an instantaneous $\mathrm{L}^{p>1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$ estimate of a form originally due to Chou & Kwong in [CK20]. Our proof has the bonus of clarifying the dependencies of the constant $C$ below.
Theorem 1.11 (Instantaneous $\mathrm{L}^{p>1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$ estimate; cf. [CK20, Proposition 2.1]).
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a Graphical Curve Shortening Flow starting from $u_{0}:=u(0)\in\mathrm{L}^{p}(J_{1})$ for some $1<p<\infty$. Then, for some universal (that is, independent of $p$ and $u$) constant $C<\infty$, there holds
(1.12) $|u(0,t)|\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right)$
@@
By combining Theorem 1.11 with Evan & Spruck’s estimate ([ES92]), one can deduce an instantaneous $\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate of the gradient of a GCSF starting from an $\mathrm{L}^{p>1}_{\mathrm{loc}}$–function. Informally, the gradient estimates assert that $\log(|\partial_{x}u(0,t)|)$ blows up like $t^{-(p+1)/(p-1)}$ for times $t\ll 1$ and grows linearity for times $t\gg 1$.

### v2

@@
A delayed interior area-to-height estimate for the Curve Shortening Flow
Arjun Sobnack [^1] Department of Mathematics, The University of Texas at Austin, TX 78712, USA. Work completed at the Warwick Mathematics Institute, University of Warwick, Coventry, CV4 7AL, UK.
Email: arjun.sobnack@austin.utexas.edu
Website: sites.google.com/view/sobnack
(31st March 2026)
?abstractname?
The principle of delayed parabolic regularity for the Curve Shortening Flow—that if two evolving curves bound a region of area $\mathcal{A}$, then, starting from time $\mathcal{A}/\pi$, the regularity of one curve is controllable in terms of the time elapsed, the area $\mathcal{A}$ and the regularity of the other curve—was proposed by Topping & the author in [ST24a], where they also provided a number of graphical situations in which their delayed regularity framework is valid. In this paper, we generalise some of the results in [ST24a] within the graphical setting, ultimately by showing that there holds an interior graphical estimate for the Curve Shortening Flow in the spirit of the proposed framework. We also provide a few applications of our estimate, such as the existence of Graphical Curve Shortening Flows starting weakly from Radon measures without point masses.
@@ 1 Introduction
Following on from Huisken’s seminal paper ([Hui84]) on the MCF of smooth, closed and mean-convex hyper-surfaces in $\mathbb{R}^{n+1}$ for $n\geq 2$, Gage & Hamilton tackled the question of the long-time behaviour of the CSF, i.e. the case of $n=1$. In [GH86] they showed that the evolution $0,T)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$ of a given convex smooth loop $\gamma_{0}=:\gamma(0)$ exists for a maximal time $T<\infty$ dictated only by the area of the region enclosed by $\gamma_{0}$. Moreover, they precisely described the behaviour of the evolution of the loop as time approaches $T$: the curve $\gamma(t)$ smoothly approaches a circle of radius $\sqrt{2(T-t)}$} in the sense of blow-ups as $t\nearrow T$. Shortly after, Grayson showed in [Gra87] that the evolution of an arbitrary smooth loop in the Euclidean plane becomes convex before encountering any singularities, at which point Gage & Hamilton’s work applies.
@@
Some of the pioneering works in this direction are [EH89, EH91] due to Ecker & Huisken. They showed that, given any [^2] Ecker & Huisken’s work considered the more general case of the MCF of graphical hyper-surfaces. graph $\gamma_{0}=\mathrm{graph}(u_{0})\subset\mathbb{R}^{2}$ of a locally Lipschitz function $u_{0}:\mathbb{R}\mapsto\mathbb{R}$, there exists a CSF $(0,\infty)\ni t\mapsto\gamma(t)=\operatorname{graph}(u(\,\cdot\,,t)\!)\subset\mathbb{R}^{2}$, defined for all time, which attains the initial datum $\gamma_{0}$ locally uniformly, and which remains graphical over the same line; we call this flow the the Ecker–Huisken flow of $u_{0}$. They achieved their existence theory by appealing directly to the Graphical Curve Shortening Flow (GCSF) equation, which the graphing function $u:\mathbb{R}\times(0,T)\mapsto\mathbb{R}$ necessarily satisfies. A key step in their proof is to establish a priori estimates for the GCSF equation of smooth graphing functions, which propagate an initial local Lipschitz bound forward to the CSF, provided one moves in a little. Such bounds are stable under locally uniform limits, and thus allowed Ecker & Huisken to construct solutions via an approximation procedure.
Around the same time, Angenent developed his Intersection Principle (see [Ang88]) and used it in [Ang91] to construct solutions to the CSF starting from Lipschitz loops in the uniform sense. As in Ecker & Huisken’s work, a key step is to establish an a priori estimate for smooth loops which propagates a Lipschitz bound forward to the flow. Angenent introduced a new and remarkably geometric technique to establish his estimates, which has found much success in other situations—see, for example, [Ilm92, CK01, Clu04, NT05, BR24], and more. In particular, Topping & the author used Angenent’s technique in [ST24a] to establish sharp delayed estimates for the GCSF, which fit into the framework to be recalled in §1.1.
@@
The running theme in the works [EH89, EH91, Ang91] above is that a strong regularity condition is shown to be preserved by the CSF. In the past three decades, significant progress has been made in the graphical situation towards instantaneous estimates, which take a weak initial regularity constraint and deduce a strong regularity condition on the CSF for strictly positive times. Evans & Spruck established in [ES92] (see also [CM04]) that an $\mathrm{L}^{\infty}_{\mathrm{loc}}$–bound on an initial graphing function induces an instantaneous $\mathrm{L}_{\mathrm{loc}}^{\infty}$–bound on the gradient of the subsequent flow, provided one moves in a little. Their result was later extended, independently, by Clutterbuck and Nagase & Tonegawa in [Clu04] and [NT05] respectively; in particular, Clutterbuck used her generalisation to establish an existence theory for a family of parabolic equations which contains graphical instances of the CSF and MCF. For $\mathrm{L}^{p}_{\mathrm{loc}}$–bounds when $1<p<\infty$, Chou & Kwong established in [CK20] an a priori estimate which takes an initial $\mathrm{L}^{p}_{\mathrm{loc}}$–bound on a graphing function instantaneously to a $\mathrm{L}^{\infty}_{\mathrm{loc}}$–bound on the gradient of the subsequent flow, provided one moves in a little; they used their estimate to establish the existence of GCSFs starting from graphs of $\mathrm{L}^{p}_{\mathrm{loc}}$–initial data. They also established a uniqueness result from $\mathrm{L}^{p}$–initial data, building on Chou & Zhu’s uniqueness theory ([CZ98]); see also [DS23].
For the non-graphical case, see, for example, the works of Angenent ([Ang90]; loosely, initial curves whose curvature is $\mathrm{L}^{p}$ for some $p\gg 1$) and Lauer ([Lau13]; finite-length Jordan curves).
One of the aims of this paper is to extend the existence theory to include, loosely, graphs of Radon measures without point masses. In some sense, the strategy we follow is familiar: we first establish an a priori estimate for smooth initial data which is stable under weak convergence of measures, and then use an approximation procedure to pass to a non-smooth limit. However, the a priori estimate we establish is pronouncedly unfamiliar in comparison to the previously-mentioned estimates in [GH86, EH89, EH91, Ang90, Ang91, ES92, Lau13] etc., as well as, to the author’s best knowledge, all other estimates (except, of course, for those in [ST24a]) from the CSF literature, which, recall, start to hold instantaneously. Our estimate is a delayed parabolic a priori estimate, and in our setting it is, in fact, impossible for the estimate to hold before a so-called magic time, which is determined completely by the ‘area’ under the initial graph. In particular, we apply our estimate slightly differently to the applications in, for example, [EH89, EH91, Ang90, Ang91, Lau13, Clu04, CK20].
@@ 1.1 A novel framework
In [ST24a] (see also [ST24b]), Topping & the author proposed a version of the following framework by which the CSF regularises a curve:
Given two smooth families $0,T)\ni t\mapsto\gamma^{-}(t)\subset\mathbb{R}^{2}$ and $0,T)\ni t\mapsto\gamma^{+}(t)\subset\mathbb{R}^{2}$ of properly embedded curves in the Euclidean plane evolving under the Curve Shortening Flow, for which $\gamma^{-}(t)$ and $\gamma^{+}(t)$ are disjoint, are either both compact or both non-compact (for all $t\in[0,T)$) and bound an evolving connected region of fixed and finite area $\mathcal{A}$, define the magic time to be $t_{\star}:=\mathcal{A}/\pi$. Then, after flowing beyond the magic time $t_{\star}$, one expects to start being able to control the regularity of $\gamma^{+}(t)$ for times $t_{\star}<t\ll T$ in terms of the time $t-t_{\star}$ elapsed, the area $\mathcal{A}$ and the regularity of $\gamma^{-}(t)$.
Several situations in which the framework above holds exactly as stated were provided in [ST24a]. Most relevant to this paper is the case where $t\mapsto\gamma^{-}(t)$ is the static flow of the xx–axis:
@@
Let $u:\mathbb{R}\times[0,\infty)\mapsto\mathbb{R}_{\geq 0}$ be a smooth positive Graphical Curve Shortening Flow starting from $u_{0}:=u(\,\cdot\,,0)\in\mathrm{L}^{1}(\mathbb{R})$. Let
$\mathcal{A}:=\|u_{0}\|_{\mathrm{L}^{1}(\mathbb{R})}=\int_{x\in\mathbb{R}}u_{0}(x)\,\mathrm{d}x$ (1.2)
@@
$t_{\star}:=\frac{\mathcal{A}}{\pi}.$ (1.3)
@@
$\|u(\,\cdot\,,t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leq\sqrt{t}\cdot H\left(\frac{t}{\mathcal{A}}\right)$ (1.4)
@@
Recall that a GCSF is a smooth function $u:J\times(0,T)\mapsto\mathbb{R}$ which satisfies the quasi-linear parabolic equation
$\partial_{t}u=\frac{1}{1+(\partial_{x}u)^{2}}\partial^{2}_{x}u=\partial_{x}[\tan^{-1}(\partial_{x}u)],$ (GCSF)
where $J\subseteq\mathbb{R}$ is an interval. A solution to (GCSF) induces a family of curves $(0,T)\ni t\mapsto\gamma(t)=\operatorname{graph}(u(\,\cdot\,,t)\!)$ which solve the CSF equation
$\left\langle\partial_{t}\gamma,\bm{n}\right\rangle=\kappa,$ (CSF)
@@
When an initial condition, say $u_{0}=:u(\,\cdot\,,0)$ for (GCSF) or $\gamma_{0}=:\gamma(0)$ for (CSF), is imposed, the regularity of the initial datum determines the mode in which the flow converges backward in time. For example, in Chou & Kwong’s work [CK20], since $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$ for some $p>1$, they insist that $u(\,\cdot\,,t)\to u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\mathbb{R})$ as $t\searrow 0$. In the case that $\gamma_{0}$ is continuous, we write $\gamma:J\times[0,T)\mapsto\mathbb{R}^{2}$ or $0,T)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$. In the case that $u_{0}$ is continuous, we write $u\in\mathrm{C}^{0}_{\mathrm{loc}}(\mathbb{R}\times[0,T)\!)$. And in the case that $u_{0}$ is smooth, we write $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times[0,T)\!)$.
In [ST24a], Topping & the author also provided the following construction, which shows that the proposed magic time $t_{\star}=\mathcal{A}/\pi$ is sharp:
@@
$\sup_{n\in\mathbb{N}}v_{n}(0,s)=\infty\qquad\text{ and }\qquad\sup_{n\in\mathbb{N}}\|v_{n}(\,\cdot\,,t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}<\infty.$ (1.7)
As applications of Theorem 1.1, Topping & the author gave a new proof of an instantaneous $\mathrm{L}^{p}$-to-$\mathrm{L}^{\infty}$–estimate ([ST24a, Theorem 1.10]), where $1<p<\infty$, and established the existence of GCSFs starting from $\mathrm{L}^{1}$–initial datum ([ST24a, Theorem 1.11]). The existence result in this paper is a generalisation of the existence result of Topping & the author, but uses a slightly different method.
@@ 1.2 Main results
Throughout this paper, we use the notation $J_{r}(a)$ to mean the interval $(a-r,a+r)\subset\mathbb{R}$ thought of as a spatial domain. We also abbreviate $J_{r}(0)$ to $J_{r}$. Any instance of a $J$ without the decoration as before will refer to an interval of unspecified length, centre or closedness. This is done purely to distinguish the spatial and temporal domains at the level of notation.
Given a spatial domain $\Omega\subseteq\mathbb{R}$, for each $k\in\mathbb{N}_{0}\cup\left\{\infty\right\}$ we denote by $\mathrm{C}^{k}(\Omega)$ the space of kk–times differentiable functions all of whose derivatives are bounded and uniformly continuously, and for each pair $(k,\alpha)\in\mathbb{N}_{0}\times(0,1]$ we denote by $\mathrm{C}^{k,\alpha}(\Omega)\subset\mathrm{C}^{k}(\Omega)$ the sub-space of functions whose $k$th derivative is $\alpha$–Hölder continuous. We denote by $\mathrm{C}^{k}_{\mathrm{loc}}(\Omega)$ and $\mathrm{C}^{k,\alpha}_{\mathrm{loc}}(\Omega)$ their respective local spaces. For a space-time domain $\mathcal{D}$, for example $\mathcal{D}=J_{1}\times(0,T)$, we denote by $\mathrm{C}^{k;l}(\mathcal{D})$ and $\mathrm{C}^{k,\alpha;l,\beta}(\mathcal{D})$ the corresponding parabolic versions of the respective spaces before, and by $\mathrm{C}^{k;l}_{\mathrm{loc}}(\mathcal{D})$ and $\mathrm{C}^{k,\alpha;l,\beta}_{\mathrm{loc}}(\mathcal{D})$ their localised counterparts. This notation is consistent with the notation introduced in §1.1 to specify the mode in which a GCSF attains it initial datum.
Given any function $f:\Omega\times\mathcal{I}\mapsto\mathbb{R}$ of space-time, where $\mathcal{I}$ is one of $(0,T)$, $[0,T)$, $(0,T]$ or $[0,T]$, we denote the tt–time slice of $f$, for each $t\in\mathcal{I}$, by $f(t):=f(\,\cdot\,,t):\Omega\mapsto\mathbb{R}$. If $\mathcal{X}$ is a function space for which $f(t)\in\mathcal{X}$ for each $t\in\mathcal{I}$, and the mapping $\mathcal{I}\ni t\mapsto f(t)\in\mathcal{X}$ is continuous, then we write $f\in\mathrm{C}^{0}_{\mathrm{loc}}(\mathcal{I};\mathcal{X})$.
In this paper, we generalise some of the results in [ST24a]—namely Theorem 1.1, the $\mathrm{L}^{p}$-to-$\mathrm{L}^{\infty}$–estimate and the $\mathrm{L}^{1}$–existence result—by demonstrating that a local, but still graphical, version of Theorem 1.1 holds. The crucial new localised estimate is
Theorem 1.8 (Delayed $\mathrm{L}^{1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate).
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a Graphical Curve Shortening Flow starting from $u_{0}:=u(0)=u(\,\cdot\,,0)\in\mathrm{L}^{1}(J_{1})$, and suppose that
$t_{\star}:=\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}=\frac{1}{\pi}\int_{x\in J_{1}}|u_{0}(x)|\,\mathrm{d}x<T.$ (1.9)
Then for all $\delta>0$, there exists a constant $C=C(\delta)<\infty$ such that the bound
$|u(0,t)|\leq C+\frac{1}{2}\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}+\frac{\pi t}{2}$ (1.10)
holds for all $t<T$ satisfying
$t\geq(1+\delta)\cdot t_{\star}.$ (1.11)
@@
For generalisations of a different flavour, the interested reader is directed to, for example, [ST26]. Topping & the author expect [ST24b] to play a role in further generalisations.
Once a height bound has been established, Evans & Spruck’s estimate ([ES92]) implies that an interior gradient bound holds an instant later, at which point standard parabolic regularity theory (see e.g. [LSU88, Kry96]) can be used to deduce that higher-order bounds hold after one more instant. For details, see [Sob24]. We shall use this argument in the proof of our existence result, which will be stated momentarily.
One advantage of the local estimate (1.10) from Theorem 1.8 is that it is applicable to CSFs that are only locally graphical. Specifically, provided one ensures that an arc of a CSF remains a graph over a fixed line segment for long enough, our estimate implies local gradient bounds, and consequently higher-order derivative bounds. In particular, one can deduce higher-order covariant derivative bounds on the geodesic curvature, which are by nature geometric as well as analytic.
@@
As an application of Theorem 1.8, in §4 we relatively swiftly deduce an instantaneous $\mathrm{L}^{p}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$ estimate for $1<p<\infty$ of a form originally due to Chou & Kwong in [CK20]. Our proof has the bonus of clarifying the dependencies of the constant $C$ below.
Theorem 1.12 (Instantaneous $\mathrm{L}^{p}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate for $p>1$; cf. [CK20, Proposition 2.1]).
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a Graphical Curve Shortening Flow starting from $u_{0}:=u(0)\in\mathrm{L}^{p}(J_{1})$ for some $1<p<\infty$. Then for some universal (that is, independent of $p$ and $u$) constant $C<\infty$ there holds
$|u(0,t)|\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right)$ (1.13)
@@
By combining Theorem 1.12 with Evan & Spruck’s estimate ([ES92]), one can deduce an instantaneous $\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate for the gradient of a GCSF starting from an $\mathrm{L}^{p}_{\mathrm{loc}}$–function, where $1<p<\infty$. Informally, the gradient estimates assert that $\log(|\partial_{x}u(0,t)|)$ blows up like $t^{-(p+1)/(p-1)}$ for times $t\ll 1$ and grows linearity for times $t\gg 1$.